Marcel Co. is growing quickly. Dividends are expected to grow at a 25 percent rate for the next 3 years, with the growth rate falling off to a constant 5 percent thereafter.

  

Required:

If the required return is 12 percent and the company just paid a $2.60 dividend. what is the current share price? (Do not round your intermediate calculations.)


rev: 09_18_2012

$62.69

$60.36

$63.97

$58.16

$65.25
Answer #1

D1=(2.6*1.25)=3.25

D2=(3.25*1.25)=4.0625

D3=(4.0625*1.25)=5.078125

Value after year 3=(D3*Growth Rate)/(Required rate-Growth rate)

=(5.078125*1.05)/(0.12-0.05)

=$76.171875

Hence current price=Future dividends and value*Present value of discounting factor(rate%,time period)

=3.25/1.12+4.0625/1.12^2+5.078125/1.12^3+$76.171875/1.12^3

=$63.97(Approx).